Peter and John are two brothers. Peter is older than John. Given that peter's age is (5x-4) years and john's age is (2x+1) years.
A) write down an expression in terms of X, for how much older Peter is than John and simplify your answer.
B) if Peter is ten years older than John, calculate the
I) value of x
II) ages of Peter and John.

Found 2 solutions by JulietG, macston:
Answer by JulietG(1812)

For (A), P = J + 10
To solve (B):
P = J + 10
(5x-4) = (2x+1) + 10
5x - 4 = 2x + 11
Subtract 2x from each side, and add 4 to each side
3x = 15
Divide by 3
x = 5
Therefore, Peter is 5(5) - 4, or 21. John is 2(5)+1, or 11.
.

Answer by macston(5194)

.
P=Peter's age; J=John's age
.
A).
P-J=(5x-4)-(2x+1)=3x-5 .
ANSWER A): Peter is 3x-5 older than John.
.
B).
10=3x-5
15=3x
5=x
.
I).
ANSWER I): Value of x is 5.
.
II).
P=5x-4=5(5)-4=25-4=21 .
ANSWER IIA). Peter is 21 years old.
J=2x+1=2(5)+1=10+1=11 .
ANSWER IIB). John is 11 years old.
.